-4x+y=-5
2x+y=7
I need to solve this by graphing, substitution or elimination

Answer by rfer(16322)

y=4x-5
2x+4x-5=7
6x=12
x=2
---------------
2*2+y=7
7=3
------------
-4*2+3=-5
-8+3=-5